OpenStudy (anonymous):

0=6x^2-10x-4 solve by factoring

OpenStudy (anonymous):

x=2 , - 1/3

OpenStudy (anonymous):

How do I solve it by factoring?

OpenStudy (anonymous):

Solve for x over the real numbers: 0 = 6 x^2-10 x-4 Reverse the equality in 0 = 6 x^2-10 x-4 in order to isolate x to the left hand side. 0 = 6 x^2-10 x-4 is equivalent to 6 x^2-10 x-4 = 0: 6 x^2-10 x-4 = 0 Factor the left hand side. The left hand side factors into a product with three terms: 2 (x-2) (3 x+1) = 0 Divide both sides by a constant to simplify the equation. Divide both sides by 2: (x-2) (3 x+1) = 0 Solve each term in the product separately. Split into two equations: x-2 = 0 or 3 x+1 = 0 Look at the first equation: Solve for x. Add 2 to both sides: x = 2 or 3 x+1 = 0 Look at the second equation: Isolate terms with x to the left hand side. Subtract 1 from both sides: x = 2 or 3 x = -1 Solve for x. Divide both sides by 3: Answer: | | x = 2 or x = -1/3

OpenStudy (anonymous):

I get that but arent you suppose to multiply 6 and negative 4 and get -24. then you would use -6 and -4 because that would equal -10 6x^2-6x+-4x-4 the gcf would be 6 and -4 x-1 and x+1 this is where I get lost I don’t understand why they aren’t the same?

OpenStudy (anonymous):

do i understand how i did it